# Robust Pronoun Use Fidelity with English LLMs: Are they Reasoning, Repeating, or Just Biased?

### v1

@@ Abstract
Robust, faithful and harm-free pronoun use for individuals is an important goal for language models as their use increases, but prior work tends to study only one or two of these components at a time. To measure progress towards the combined goal, we introduce the task of pronoun use fidelity: given a context introducing a co-referring entity and pronoun, the task is to reuse the correct pronoun later, independent of potential distractors. We present a carefully-designed dataset of over 5 million instances to evaluate pronoun use fidelity in English, and we use it to evaluate 37 popular large language models across architectures (encoder-only, decoder-only and encoder-decoder) and scales (11M-70B parameters). We find that while models can mostly faithfully reuse previously-specified pronouns in the presence of no distractors, they are significantly worse at processing she/her/her, singular they and neopronouns. Additionally, models are not robustly faithful to pronouns, as they are easily distracted. With even one additional sentence containing a distractor pronoun, accuracy drops on average by 34%. With 5 distractor sentences, accuracy drops by 52% for decoder-only models and 13% for encoder-only models. We show that widely-used large language models are still brittle, with large gaps in reasoning and in processing different pronouns in a setting that is very simple for humans, and we encourage researchers in bias and reasoning to bridge them.
@@ 1 Introduction
Figure 1: We evaluate model accuracy at using the correct pronoun for an entity when provided with an explicit introduction and 0-5 distractor sentences. Llama-2-70B and RoBERTa-large show large accuracy drops with just one distractor.
@@
Accordingly, language technology should acknowledge the sensitivity of faithful and correct pronoun use. To this end, many studies have explored how large language models handle pronouns, showing that they stereotypically associate pronouns and occupations (Kurita et al., 2019), reason about co-referring pronouns and entities better when they conform to stereotypes (Tal et al., 2022), fail when exposed to novel pronoun phenomena such as neopronouns (Lauscher et al., 2023), and cannot consistently reuse neopronouns during generation (Ovalle et al., 2023a). These shortcomings create quality of service differences and cause representational harm, amplifying discrimination against certain pronoun users (Blodgett et al., 2020; Dev et al., 2021).
However, (i) studies on stereotypical associations tend to test for intrinsic bias, with brittle methodology (Seshadri et al., 2022) that may not translate to extrinsic bias or harms in downstream applications (Goldfarb-Tarrant et al., 2021), (ii) studies on reasoning with pronouns tend to focus on coreference resolution, the ability to reason about connections between pronouns and entities, which may not translate to a model’s use of pronouns, and (iii) studies on pronoun use tend to examine simplistic pronoun reuse with no more than one entity. A key research question that has gone unanswered thus far is: How robustly faithful are models to one’s pronouns? We address this research gap by presenting the first study measuring the robustness and fidelity of models’ pronoun use.
@@
Contributions.(1) To investigate whether models can reason robustly about pronouns in context, we propose a new task, pronoun use fidelity: given a context in which a co-referring entity and pronoun are introduced, the task is to reconstruct the pronoun later, independent of potential distractors. (2) We present a novel, large-scale dataset consisting of 7,000 up to 2 million instances depending on the setup, carefully designed to evaluate this task. (3) We study pronoun use fidelity across 37 popular language models covering a range of architectures and scales, and we analyze model errors to examine whether models are reasoning, repeating, or just biased.
@@
First, we examine pronoun predictions in the absence of context as previous studies have done, to establish a bias baseline and examine the consistency of model predictions across case (§5). Next, we evaluate whether models can overcome their biased pronoun predictions by showing them an introductory context to explicitly establish what pronoun to use (§6). We find that all models are good at this task, and bigger models are better. Then, we attempt to distract models by inserting additional sentences using a different pronoun to talk about another person (§7). Models are not robust to distractors, and even one distractor sentence vastly deteriorates model performance as shown in Figure 1. This result holds even at the scale of 70 billion parameter models, despite how simple this task is for humans. Finally, in a detailed error analysis (§8), we examine the extent to which model errors can be attributed to distraction or falling back to intrinsic bias. Overall, we show a large gap in robust model reasoning about pronouns in a simple setting, and we find that encoder-only and decoder-only models behave in fundamentally different ways, not only in their performance on this task, but also in the types of errors they make. We release all code and data at https://github.com/uds-lsv/pronoun-use-fidelity.
2 Pronoun Use Fidelity Task
Given a context in which a co-referring entity and pronoun are introduced, the task is to reconstruct the pronoun later in a sentence about the entity, independent of potential distractors.
@@ 2 Pronoun Use Fidelity Task
3 Data
We begin by discussing our dataset construction procedure to evaluate pronoun use fidelity at scale. It consists of a three-step pipeline allowing us to create naturalistic instances of our task while ensuring a correct answer: template creation (§3.1), template assembly (§3.2), and validation (§3.3).
Our data covers 60 occupations (see Appendix A) and four third-person pronouns (he, she, they and xe) in three cases (nominative, accusative and possessive dependent). Our occupations (and participants for the distractor sentences) are chosen from Winogender schemas (Rudinger et al., 2018), as their bias characteristics are well-studied in NLP fairness literature. In addition to the English standard masculine (he/him/his) and femininine (she/her/her) pronouns, we heed Lauscher et al.’s (2022) call for more inclusive NLP research by examining two more pronoun sets that are less well-studied in the fairness literature: singular they (they/them/their), the pronoun of choice of over 75% of respondents to the Gender Census (Lodge, 2023), and xe/xem/xyr, the most popular neopronoun according to the same census.
@@ 3.1 Template creation
Task templates. We create one task sentence template per occupation and pronoun case, for a total of 180 templates, designed to create an unambiguous coreference with the occupation only. For instance, charges for preparing tax returns can only belong to an accountant, never a taxpayer, which is the corresponding participant.
Context templates. Semantically bleached templates such as He is an accountant are well-established in the literature for testing word embedding associations (Caliskan et al., 2017), but it is unnatural to use more than one consecutively. On the other hand, natural corpora like Levy et al. (2021) contain occupation-specific sentences that cannot be used more generally. As a compromise between a more controlled setting and naturalistic data, we create templates with generic themes that apply to all humans, e.g., universal emotions and sensations (hungry/full, tired/energetic, unhappy/happy, etc.). We create 10 context templates per pronoun case, for a total of 30 templates.
@@ 3.2 Template assembly
We then combine and instantiate the previously created templates to assemble our data instances. First, we select a task sentence for a certain occupation and pick one of four pronouns in the correct case to use as the ground truth label for the task. The occupation becomes entity $e_{a}$ and the pronoun is $p_{a}$. We then create an introduction sentence by sampling one of the 10 context templates with the matching case, and instantiating it with the selected occupation $e_{a}$ and pronoun $p_{a}$. The simplest version of the pronoun use fidelity task includes just this introduction sentence followed by the task sentence. Instantiating 10 introductory templates with 4 different pronouns and pairing them with task templates for 60 occupations across 3 cases gives us a total of 7,200 unique instances for this version of the task.
To create more complex data instances, we insert a variable number of distractor sentences between the introduction and task sentences. The first distractor sentence always explicitly establishes a coreference between the participant and the distractor pronoun, and any subsequent distractor sentences implicitly continue this coreference by using just the pronoun. We first select a distractor pronoun $p_{b}$ from the three unused pronouns. We sample templates from the remaining context templates [^1] For more details on this sampling, refer to Appendix B., instantiating them with this pronoun and a distractor entity $e_{b}$, the participant that the occupation is paired with. For the one-distractor case, instantiating 4 templates with 3 unused pronouns gives us 86,400 unique instances.
Data type Number of instances
@@
Table 1: Number of dataset instances. Pronoun use fidelity instances consist of task instances combined with introductory contexts and optional distractors. We subsample 3 sets of 2,160 sentences (of the total number of instances we created).
Our stackable dataset design allows us to generate a vast amount of data of varying lengths, which we believe reflects varying levels of difficulty for models. We subsample these datasets with three random seeds for the rest of our evaluation, ensuring that all occupations, cases, pronoun declensions and distractor pronouns are equally represented in each subsampled set of 2,160 sentences. Data statistics are shown in Table 1.
3.3 Data validation
To verify that the pronoun use fidelity task is valid and easy for humans, we validate a subset of the combined sentences, in addition to validating all the individual sentence templates. All annotator instructions are provided in Appendix C.
Sentence templates. Two authors with linguistic training iteratively created and validated the sentence templates for grammaticality and coreference correctness until we reached consensus. These were independently verified by an annotator who also rated 100% of the sentences as grammatical and with the correct coreferences.
Pronoun use fidelity task. We sampled 100 instances with each possible number of distractors (0-5), for a total of 600 instances. One author and one annotator had to fill in the pronoun and they each performed with 99.8% accuracy. [^2] They disagreed on non-overlapping instances which appeared to be random slips.
@@ 4 Experimental Setup
We first list the models we experiment with, and explain our evaluation methods and metrics. Further details are provided in Appendix D.
@@ 4.1 Models
We experiment with a total of 37 transformer-based language models, chosen to evaluate the effects of architecture and scaling. Table 2 shows all models we experiment with. Our 9 encoder-only models are from the BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), ALBERT-v2 (Lan et al., 2020) and MosaicBERT (Portes et al., 2023) model families, as the first three remain well-used in NLP fairness literature, and the last is trained on much more data. As for our 20 decoder-only models, we select the popular Llama-2 (Touvron et al., 2023) model family, as well as OPT (Zhang et al., 2022) and Pythia (Biderman et al., 2023) for their large range of model sizes. We also experiment with eight popular chat models that are further trained with instruction-tuning and reinforcement learning, to evaluate their instruction-following behaviour on this task; specifically, we use decoder-only Llama-2-chat models (Touvron et al., 2023) and encoder-decoder FLAN-T5 models (Chung et al., 2022).
@@ 4.2 Obtaining predictions
Decoder-only models. For the majority of our experiments, we follow Hu and Levy (2023) in taking direct measurements of probabilities as a proxy for models’ metalinguistic judgements. We verbalize four versions of each data instance, i.e., we fill in the blank with each of the four pronouns we consider, creating four options. We then compute sentence-level model log likelihoods for these options, and select the option with the highest log likelihood as the model’s choice.
Encoder-only models. For comparable evaluation across model architectures and pronouns, we use pseudo log likelihoods for encoder-only models (Salazar et al., 2020; Kauf and Ivanova, 2023). We do not use masked token prediction due to tokenization issues with neopronouns (Ovalle et al., 2023b); briefly, we want xe to be tokenized as it normally would be (which is often as two tokens) rather than as a single UNK token.
Chat models. In line with their most naturalistic usage scenarios, we evaluate chat models (FLAN-T5 and Llama-2-chat) separately using prompting. We report the range of performance with 10 different prompts using boxplots, following Sclar et al. (2024). Our prompts, shown in Appendix E, were inspired by the prompts to elicit coreferences in the FLAN collection (Longpre et al., 2023).
@@ 4.3 Metrics
On the task sentences with no context, as there is no correct answer, we calculate the Pearson correlation coefficient between each model’s nominative, accusative and possessive pronoun predictions to evaluate the models’ case consistency.
On the pronoun fidelity task, as there is a correct answer, we compute accuracy. We report mean accuracy over the three randomly sampled subsets of our dataset, and show the standard deviation with error bars or shading. Where possible, we perform significance testing with a Welch’s t-test and a threshold of 0.05. We use human performance as our ceiling, and compare models evaluated with (pseudo) log likelihoods to a baseline of randomly selecting 1 of the 4 pronouns (i.e., 25%).
@@ 5 Model Predictions with No Context
Model pronoun prediction with no context is the task studied in numerous papers on pronominal gender bias (Kurita et al., 2019, inter alia), which show associations between occupations and pronouns based on social gender stereotypes, e.g., doctor-he and nurse-she. However, most of them ignore cases beyond the nominative (Munro and Morrison, 2020). If model associations between occupations and pronouns are primarily explained by social bias and occupational statistics, then we should see correlations between pronoun predictions across cases. In other words, if a model associates accountants with he, then it might also associate accountants with him and his. Therefore, we first examine model pronoun predictions on our task sentences with no context, and evaluate the case consistency of pronoun predictions.
Figure 2: Distribution of Pearson correlations of per-occupation pronoun predictions of different cases: nom(inative), acc(usative) and poss(essive).
Figure 3: Pronoun use fidelity of all models with an introductory context and no distractors. Accuracy is averaged across all pronouns and cases, and is above chance (0.25) but below human performance (1.0).
@@
she/her/her Refer to caption
they/them/their Refer to caption
xe/xem/xyr Refer to caption
Figure 4: Pronoun use fidelity with an introductory context and no distractors, split by pronoun series.
Figure 2 shows the distribution of Pearson correlations between per-occupation pronoun predictions of different cases (individual model correlations are provided in Appendix F). The range is large, e.g., --0.46 for RoBERTa-base compared to 0.47 for Pythia-160M for nominative-accusative prediction correlations, showing that there is a lot of variation in pronoun associations across case and models. Only nominative and possessive pronoun predictions are positively correlated for all but one model (ALBERT-base-v2), and accusative pronouns seem to behave quite differently; Llama-2-13B, for instance, has the highest nominative-possessive correlation of all models, 0.55, but nominative-accusative and accusative-possessive correlations of 0.00 and --0.01 respectively. As pronoun predictions seem not to line up well across case, these results suggest that biased pronoun predictions in this context just mean statistically biased predictions, adding to the existing evidence that templates are a brittle method of evaluating social biases in models (Selvam et al., 2023; Seshadri et al., 2022). Next, we examine whether models can override their intrinsic statistical biases on these templates when provided with a pronoun to use.

### v2

# Robust Pronoun Fidelity with English LLMs: Are they Reasoning, Repeating, or Just Biased?

@@ Abstract
Robust, faithful and harm-free pronoun use for individuals is an important goal for language models as their use increases, but prior work tends to study only one or two of these characteristics at a time. To measure progress towards the combined goal, we introduce the task of pronoun fidelity: given a context introducing a co-referring entity and pronoun, the task is to reuse the correct pronoun later. We present RUFF, a carefully-designed dataset of over 5 million instances to measure robust pronoun fidelity in English, and we evaluate 37 popular large language models across architectures (encoder-only, decoder-only and encoder-decoder) and scales (11M-70B parameters). When an individual is introduced with a pronoun, models can mostly faithfully reuse this pronoun in the next sentence, but they are significantly worse with she/her/her, singular they and neopronouns. Moreover, models are easily distracted by non-adversarial sentences discussing other people; even one additional sentence with a distractor pronoun causes accuracy to drop on average by 34%. Our results show that pronoun fidelity is neither robust, nor due to reasoning, in a simple, naturalistic setting where humans achieve nearly 100% accuracy. We encourage researchers to bridge the gaps we find and to carefully evaluate reasoning in settings where superficial repetition might inflate perceptions of model performance.
@@ 1 Introduction
Accordingly, it is important for large language models (LLMs) to use pronouns faithfully and without causing harm. To this end, many studies have explored how LLMs handle pronouns, showing that they stereotypically associate pronouns and occupations (Kurita et al., 2019), reason about co-referring pronouns and entities better when they conform to stereotypes (Tal et al., 2022), fail when exposed to novel pronoun phenomena such as neopronouns (Lauscher et al., 2023), and cannot consistently reuse neopronouns during generation (Ovalle et al., 2023a). These shortcomings create quality of service differences and cause representational harm, amplifying discrimination against certain pronoun users (Blodgett et al., 2020; Dev et al., 2021).
@@
However, a question that has gone unexamined thus far is: How robust is model faithfulness to pronouns when discussing more than one person? To answer this question, we propose pronoun fidelity2), a new task to investigate realistic model reasoning about pronouns, and we introduce RUFF3), a novel, large-scale dataset of over 5 million instances, carefully designed to evaluate this task. With this dataset, we present an analysis of pronoun fidelity across 37 popular language models covering a range of architectures and scales (11M-70B), to investigate whether models are reasoning, repeating, or just biased.
@@
First, we collect model pronoun predictions for occupations in the absence of context, to establish a “bias baseline” (§5). Next, we evaluate whether models can overcome their biased pronoun predictions when explicitly shown what pronoun to use in context (§6). All models are good at this task, but there are significant disparities across pronoun sets. We then test the robustness of this result by inserting naturalistic distractor sentences using a different pronoun to talk about another person (§7). Even one non-adversarial distractor sentence vastly deteriorates model performance as shown in Figure 1. Finally, in a detailed error analysis (§8), we disentangle whether model errors can be attributed to distraction or falling back to bias, finding that encoder-only and decoder-only models behave in fundamentally different ways.
Figure 1: We evaluate model accuracy at using the correct pronoun for an entity when provided with an explicit introduction and 0-5 non-adversarial distractor sentences. Llama-2-70B and RoBERTa-large show large accuracy drops with just one distractor. Accuracy is averaged over 3 data splits; standard deviation is shown with shading.
Overall, our results show that models struggle to reason about pronouns in a simple, naturalistic setting and highlight the need for careful task design to ensure that superficial repetition does not lead to inflated claims about model reasoning. We release all code and data to encourage researchers to bridge the gaps we find: https://github.com/uds-lsv/robust-pronoun-fidelity
2 Pronoun Fidelity Task
Discussing multiple individuals is natural, frequent and well-studied in discourse; we use both definite references and pronouns in natural language to establish continuity and coherence (Grosz et al., 1995). We formalize a version of these phenomena in our task: given a context in which a co-referring entity and pronoun are introduced, the task is to reconstruct the pronoun later in a sentence about the entity, independent of a limited number of potential distractors.
@@ 2 Pronoun Fidelity Task
3 RUFF Dataset
RUFF, our evaluation dataset for Robust pronoUn Fidelity at scale, is constructed with a 3-step pipeline to create naturalistic instances of our task with unambiguous answers: template creation (§3.1), assembly (§3.2), and validation (§3.3).
Our data covers 60 occupations (see Appendix B) and four third-person pronouns (he, she, they and xe) in three grammatical cases (nominative, accusative and possessive dependent). Our occupations (and corresponding participants for distractor sentences) are chosen from Winogender schemas (Rudinger et al., 2018), as their bias characteristics are well-studied in NLP. In addition to the English masculine (he/him/his) and feminine (she/her/her) pronouns, we heed Lauscher et al.’s (2022) call for more inclusive NLP research by examining two more pronoun sets that are less well-studied in NLP: singular they (they/them/their), the pronoun of choice of over 75% of respondents to the Gender Census (Lodge, 2023), and xe/xem/xyr, the most popular neopronoun according to the same census.
@@ 3.1 Template creation
Task templates. We create one task sentence template per occupation and grammatical case for a total of 180 templates, designed to create an unambiguous coreference with the occupation only. For instance, charges for preparing tax returns can only belong to an accountant, never a taxpayer, which is the corresponding participant.
Context templates. We create explicit (definite reference + pronoun) and implicit (pronoun-only) versions of 10 context templates per grammatical case, for a total of 30 templates. Each explicit context template begins with an entity and introduces the pronoun in a clause, e.g., The taxpayer needed coffee because their day had started very early, while implicit templates are simple sentences like Their sleep had been fitful. All templates contain generic themes, e.g., universal human emotions and sensations (hungry/full, tired/energetic, unhappy/happy, etc.). These templates can thus be applied to all occupations and participants, and allow for a controlled but still coherent setting.
@@ 3.2 Template assembly
We then instantiate and combine the previously created templates to assemble our data instances. First, we select an occupation ($e_{a}$) and one of its task templates. We pick a pronoun ($p_{a}$) to use as ground truth and instantiate a random context template with the selected occupation and pronoun. The simplest version of the pronoun fidelity task includes just this introduction sentence followed by the task sentence. Instantiating 10 templates with 4 different pronoun sets and pairing them with task templates for 60 occupations across 3 grammatical cases gives us a total of 7,200 unique instances for this version of the task.
To create more complex data instances, we insert a variable number of distractor sentences between the introduction and task sentences, discussing a participant $e_{b}$ who uses a different pronoun $p_{b}$. These are also sampled from the set of context templates (see Appendix C for details). Instantiating 4 templates with 3 unused pronouns gives 86,400 unique instances with one distractor.
Data typespace spaceNumber of instances
@@
Table 1: Number of dataset instances. Pronoun fidelity instances consist of task instances combined with introductory contexts and optional distractors. We subsample 3 sets of 2,160 sentences (of the total number of instances we created).
Our stackable dataset design allows us to generate a vast amount of data of varying lengths, giving us a controlled setting to evaluate context effects on model predictions. We subsample the data with three random seeds for the rest of our evaluation, ensuring that all occupations, cases, pronoun declensions and distractor pronouns are equally represented in each subsampled set of 2,160 sentences. All data statistics are shown in Table 1.
@@
3.3 Data validation
We validate all task and context templates. To verify that the pronoun fidelity task is easy and unambiguous for humans, we also validate a subset of these instances with 0-5 distractors. Annotator information is shown in Appendix D and all annotator instructions are provided in Appendix E.
Templates. Two authors with linguistic training iteratively created and validated sentence templates for grammaticality and correct coreferences until consensus was reached. An additional annotator independently rated 100% of the sentences as grammatical and with the correct coreferences.
Pronoun fidelity task. We sampled 100 instances with each possible number of distractors (0-5), for a total of 600 instances. One author and one annotator had to fill in the pronoun and they each performed with 99.8% accuracy. [^1] They disagreed on non-overlapping instances which appeared to be random slips.
@@ 4 Experimental Setup
We list our models, evaluation methods, and metrics. Further details are provided in Appendix F.
@@ 4.1 Models
We experiment with 37 transformer-based language models (see Table 2), chosen to evaluate the effects of architecture and scaling. Our 9 encoder-only models are from the BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), ALBERT-v2 (Lan et al., 2020) and MosaicBERT (Portes et al., 2023) model families, as the first three remain well-used in NLP, and the last is trained on much more data. As for our 20 decoder-only models, we select the popular Llama-2 (Touvron et al., 2023) model family, as well as OPT (Zhang et al., 2022) and Pythia (Biderman et al., 2023) for their large range of model sizes. In Appendix I, we also experiment with eight popular chat models that are further trained with instruction-tuning and reinforcement learning, to evaluate task performance with prompting; specifically, we use decoder-only Llama-2-chat models (Touvron et al., 2023) and encoder-decoder FLAN-T5 models (Chung et al., 2022).
@@ 4.2 Obtaining predictions
Decoder-only models. For the majority of our experiments, we follow Hu and Levy (2023) in taking direct measurements of probabilities as a proxy for models’ metalinguistic judgements. We verbalize four versions of each data instance, i.e., we fill in the blank with each of the four pronouns we consider, creating four options. We then compute instance-level model log likelihoods for these options, and select the option with the highest log likelihood as the model’s choice.
Encoder-only models. For comparability with decoder-only models, we follow the same procedure for encoder-only models, but use pseudo log likelihoods (Salazar et al., 2020; Kauf and Ivanova, 2023). We do not use masked token prediction due to tokenization issues with neopronouns (Ovalle et al., 2023b); briefly, we want xe to be tokenized “normally” (which is often as two tokens) rather than a single UNK token.
Chat models. Following common practice, we evaluate chat models (FLAN-T5 and Llama-2-chat) using vanilla and chain-of-thought prompting in Appendix I. Following Sclar et al. (2024), we show the range of expected performance with 10 different prompts (see Appendix G), inspired by the prompts to elicit coreferences in the FLAN collection (Longpre et al., 2023).
@@ 4.3 Metrics
As every instance of the pronoun fidelity task has a unique correct answer, we report accuracy averaged over the three randomly sampled subsets of our dataset. We show the standard deviation with error bars or shading. Where possible, we perform significance testing with a Welch’s t-test and a threshold of 0.05. We use human performance as our ceiling, and compare models to a baseline of randomly selecting 1 of the 4 pronouns (i.e., 25%).
Figure 2: Counts of pronoun predictions from all models, in the absence of context. Error bars indicate standard deviation across models.
@@ 5 Model Predictions with No Context
We begin by creating a “bias baseline,” i.e., obtaining pronoun predictions from models on our task sentences in the absence of any context. In Section 6, we will examine whether models can overcome this bias with reasoning when provided with context establishing a single correct answer.
[…]
As we cannot evaluate accuracy on a task with no single correct answer, we show the counts of model predictions of different pronoun declensions in Figure 2, averaged over all models. Model-specific counts are shown in Appendix H. Even though our task sentences are designed such that any pronoun set can be used grammatically, all models tend to assign higher probability to he/him/his than other pronoun sets.
Obtaining pronoun predictions without context is a popular method to measure model bias, with numerous papers (Kurita et al., 2019, inter alia) showing that associations between occupations and pronouns are based on social gender stereotypes, e.g., doctor-he and nurse-she. However, model pronoun predictions might reflect dataset artifacts such as the choice of occupations, or be a statistical accident of the chosen templates (Seshadri et al., 2022). In addition, intrinsic biases may not correlate with actual pronoun use with context (Goldfarb-Tarrant et al., 2021). In order to test for such extrinsic behaviours, the rest of this paper examines whether models can override their intrinsic statistical biases on these same templates when provided with the right pronoun to use.
Figure 3: Pronoun fidelity by model with an introductory context. Accuracy is averaged across occupations, pronouns and grammatical cases, and is above chance (0.25) but below human performance (1.0).
@@
she/her/her
(* < he/him/his)
they/them/their
(* < she/her/her)
xe/xem/xyr
(* < she/her/her)
Figure 4: Pronoun fidelity by model with an introductory context, split by pronoun series. Model accuracy is compared to chance (0.25) and human performance (1.0) and * denotes statistical significance.